# Effective programming of a photonic processor with complex interferometric structure

### v1

@@ Abstract
Reconfigurable photonics have rapidly become an invaluable tool for information processing. Light-based computing accelerators are promising for boosting neural network learning and inference [1, 2, 3] and optical interconnects are foreseen as a solution to the information transfer bottleneck in high-performance computing [4, 5, 6]. In this study, we demonstrate the successful programming of a transformation implemented using a reconfigurable photonic circuit with a non-conventional architecture. The core of most photonic processors is an MZI-based architecture [7, 8] that establishes an analytical connection between the controllable parameters and circuit transformation. However, several architectures [9, 10] that are substantially more difficult to program have improved robustness to fabrication defects. We use two algorithms that rely on different initial datasets to reconstruct the circuit model of a complex interferometer, and then program the required unitary transformation. Both methods performed accurate circuit programming with an average fidelity above 98%. Our results provide a strong foundation for the introduction of non-conventional interferometric architectures for photonic information processing.
@@ I Introduction
Integrated photonics have already significantly impacted in various areas of science and technology, including communication [11, 12, 13], quantum and classical computing [14], optical neural networks [15, 2, 3], and life sciences [16, 17]. The key feature of integrated photonic systems that drives the growth of interest is the ability to route and manipulate light states using on-chip phase modulation. Information processing tasks use interferometric waveguide circuits that enable programmable encoding of information processing algorithms using light. The benefits are the larger operational bandwidth of photonic devices and a better energy efficiency per operation [18, 19, 20].
In this study, we target photonic device programming for information processing. A typical photonic chip for this purpose includes multiple power splitters and phase shifters comprising a large multiarm interferometer [21, 22, 23]. Once a photonic processor has been fabricated, it must be properly calibrated to implement a unitary transformation on demand. This process is challenging if the interferometric circuit architecture includes complex elements, such as multiport power-splitting circuits [24, 2, 3].
@@
Common photonic architectures rely on a Mach-Zehnder interferometer (MZI) as a building block [7, 8, 25, 26]. This class of architectures bases on the well-known isomorphism of the MZI transfer matrix and the arbitrary unitary $SU(2)$ rotation. Within MZI-based architectures, the calibration can be performed for each interferometer one after another, as it is possible to isolate the effect of each individual block [21, 15, 27, 28, 29, 19].
However, this strategy is not applicable to interferometers with other types of building blocks. For example, using multiport beamsplitters [10] or beamsplitter meshes [30] turns out to be beneficial in minimizing the effects of fabrication imperfections. Interferometers of this kind have already been demonstrated [9, 31, 32] but programming a desired unitary transformation remains an unresolved challenge. The well-known issue of the crosstalk between multiple tunable phase shifters complicates the unitary programming problem even more by rendering easy calibration procedures inaccurate even in simple architectures such as [7, 8].
@@
However, there exists substantial practical interest in interferometers with $n\times n$ waveguide splitters [36, 37, 38], as they can be used as a powerful tool for photonic convolutional networks and random matrix generation [2, 39]. We will call the interferometers consisting of such couplers multiport interferometers. There is no straightforward approach to enable programming unitary transfer matrices of the interferometer with $n\times n$ splitters.
It is worth mentioning that many of the existing studies on MZI-based interferometer programming have not used the actual matrix fidelity measure to demonstrate the ability to achieve the desired unitary matrix set on a chip. Instead, in these studies, another metric called amplitude fidelity was used [28, 19, 34, 23, 35], which is inherently insensitive to the angles of the unitary transformation that are crucial for the evolution of multiphoton states through linear optical interferometers. However, amplitude fidelity has been argued to contain nearly the same amount of information, demonstrating the universal capabilities of the interferometers under study [29].
Several recent theoretical proposals demonstrated how to program an arbitrary complex interferometer in two steps [40, 41]. In the first step, the digital model of the interferometer is inferred from the experimental data set, and the digital model afterwards serves as a reliable tool to determine the phase configuration corresponding to the required unitary transfer matrix. Both methods require precise control of the phase shifts $P(\vec{\varphi})$ that in turn demand fine calibration $\vec{\varphi}=\vec{\varphi}(\vec{X})$ of the phase modulators.
@@ II.1 Optical Chip Design
Our photonic chip was manufactured using femtosecond laser writing (FSLW) in a fused silica glass sample [42]. The description of the waveguide fabrication regime can be found in Appendix A. The single-mode waveguide structure of the chip contained a target $4\times 4$ multiport interferometer and a preparation interferometer. The preparation interferometer is a mesh of three MZIs which can set an arbitrary amplitude distribution at the output of the preparation part. We use the preparation interferometer as a tool for calibrating the target interferometer and probing its transfer matrix. Both the preparation and the target interferometers are equipped with tunable thermo-optical phase shifters, which are also fabricated using the femtosecond laser writing setup.
@@
Calibration of the MZI-based interferometer is a known procedure that is performed by individually calibrating each of its constituent MZIs [21, 27, 28, 19]. We used the same principles to calibrate the preparation interferometer. More details on the optical chip design and preparation interferometer calibration can be found in Appendices B and C.
@@ II.2 Target interferometer calibration
However, in most real optical chips, even those fabricated using the most advanced technology [28], there is often an effect of mutual cross-influence between different phase shifters. This can occur because of thermal or electrical crosstalk. Due to the crosstalk effect, the optical power $P_{m}$ from the $m$-th output port of the chip can no longer be described by 2 because the approximation constants $A_{m}$ and $B_{m}$ are functions of $x_{j}$, as well as of the indices of the input and output ports of the interferometer. This leads to a complex and non-periodic dependence $P_{m}(x_{j})$. Figure 2b shows the typical dependency $P_{m}(x_{j})$ in the presence of crosstalk.
@@
where $x_{j}$ is the current applied to $j$-th heater, $\varphi^{(i)}_{0}$ is a constant bias phase shift that exists in the particular interferometer arm, and $\alpha_{ij}$ is a constant that determines the strength of the effect of the $j$-th heater on the $i$-th phase shifter. The relation between the phases and the currents flowing through the heaters can be rewritten in matrix form as follows:
@@
Moreover, any experimentally measured optical power $P_{km}$ from the applied current for any input $k$ and output $m$ ports of the interferometer and for any heater $j$ must obey (5) with the same interferometer parameters $M_{1,2}$, $\overrightarrow{\Phi_{0}}$ and $A$, as they are all generated by the same photonic chip, described by its unitary matrix $U$. Once all possible optical powers from the applied current dependencies, $P_{km}(x_{j})$, have been experimentally measured for each $k$, $m$, and $j$, they all can be simultaneously approximated using a single interferometer model, determining the optical chips’ constituents $M_{1,2}$ and the necessary calibration relations between all currents and phase shifts $\overrightarrow{\Phi_{0}}$ and $A$. The key points of the approximation over the whole dataset are given in the next subsection.
@@ II.3 The target interferometer model
where $\vec{t}_{1,2}=\{t_{1},\dots,t_{N}\}$ are two sets of real parameters that determine $M_{1,2}$, $P(\vec{\varphi}(\vec{x},\vec{\varphi}_{0},A))=diag(e^{i\varphi_{1}},e^{i\varphi_{2}},e^{i\varphi_{3}},1)$ is a diagonal matrix of the phase shifts (see Fig. 2c). In turn, the phase shifts depend on the currents $x_{j}$ and the calibration parameters $\varphi^{(0)}_{j}$ and $\alpha_{ij}$:
@@
For mode-mixing blocks $M_{1,2}$ we used the well-known unitary parameterization [7] that for the $4\times 4$ port interferometer requires nine real parameters varying from 0 to $2\pi$ each (see Appendix D for details). The total number of real parameters for our photonic chip then equals to $30$, including $9\times 2=18$ parameters for $M_{1,2}$, $3$ bias phases $\varphi^{(0)}_{j}$, and $9$ crosstalk constants $\alpha_{ij}$.
@@ III Experimental results
We gathered the data set $P_{km}(x_{j})$ and ran the approximation which yielded $M_{1}$ and $M_{2}$, $\overrightarrow{\Phi_{0}}$, and $A$, which are illustrated in Fig. 2d,e. Although no additional constraints were set for the parameterization of $M_{1}$ and $M_{2}$ they appeared to be almost equal $M_{1}\approx M_{2}$ as shown in Fig. 2d, which reflects the fact that physically in our optical chip both $M_{1,2}$ were realized as $4\times 4$ directional couplers with identical geometrical parameters (i.e. waveguide shapes, distances between waveguides and laser exposure parameters). Moreover, from Fig. 2e we can see that the bias phases satisfy $\varphi^{0}_{1}\approx\varphi^{0}_{4}$ and $\varphi^{0}_{2}\approx\varphi^{0}_{3}$, which can also be explained by the actual geometry of the phase shift layer $P(\varphi)$ on our chip (see Fig. 1).
In the phase shift layer of the target interferometer, the first and fourth waveguides have the same length, while the second and third waveguides also have the same length but are shorter than the first and fourth. This results in non-zero bias phase shifts, as shown in Fig. 2e.
The reconstructed elements of the crosstalk matrix $A$ (see Fig. 2e) correlate with our assumption about their mutual relation based on the geometry of the optical structure. The negative values of some of the elements are also predicted by our simple model. The qualitative estimation of the relative magnitudes of $a_{ij}$ can be made using the linearity of the heat equation, the assumption of logarithmic temperature decay with distance from the heater and the proportionality between the phase and temperature differences $\varphi_{i}-\varphi_{j}\sim T_{i}-T_{j}$. Assuming (endless) parallel straight heaters with a width of $r_{0}$ right above (endless) parallel straight waveguides are equally separated by $d$ from each other (see Fig. 3), one can derive the following form of the crosstalk matrix $A$:
@@ III.1 Testing the digital model of the chip
To verify the validity of our digital interferometer model, we perform the following procedure. We sampled a set of 100 random current configurations $\{x^{(n)}_{1},x^{(n)}_{2},x^{(n)}_{3}\}$ ($n=\overline{1,100}$) in the range of 0 to 15 mA and applied them to our photonic chip. For each set of currents, the chips’ unitary transformation $U^{(n)}_{exp}$ was measured directly using coherent CW radiation [43, 44] (see Appendix E for details). On the other hand, $U^{(n)}$ can be simulated from the digital chip model (6) and compared to the measured one by the matrix fidelity measure:
@@
III.2 Optical port-to-port permutations in a wide range
@@
The procedure described above for the experimental calibration of the photonic chip (and programming the chip at the same time) requires simply injecting coherent radiation into each mode of the chip, applying current to a specific heater, and recording the optical power from all output ports. After the successful approximation of the entire data set of the calibration data collected, a complete digital model of the chip is obtained. In this procedure, the wavelength $\lambda$ determines the resulting unitary transformation $U=U(\lambda)$ due to the spectral features of the mode-mixing blocks and the spectral dependence of the phase shifts. The reconstruction of the digital model can be repeated with different wavelengths $\lambda_{new}$ to retrieve the chip model $U=U(\lambda_{new})$. The multiport interferometer architecture [10] was theoretically proven to be robust to the deviations of the mode-mixing block transfer matrices. By implementing the same operation, the optical port-to-port switching, using interferometers with different mode-mixing blocks we highlight the robustness feature of the multiport architecture.
@@
To demonstrate the wide spectral range of our photonic chip capabilities and its potential to perform predefined optical transformations with practical applications, we have experimentally performed optical switching at seven different wavelengths between $915$ and $975$ nanometers. The optical switching is a particularly demanding task for an interferometric circuit, especially based on MZI blocks, because it requires precise calibration of the power couplers and phase shifters. We first calibrated our optical chip for seven wavelengths using a continuously tunable CW laser (Toptica CTL 950), which has a range of wavelengths between $\lambda_{m}=915$ and $975$. We then obtained digital models $U(\vec{x},\lambda_{m})$ for each of these wavelengths. Then, using the known digital model of the chip, $U(\vec{x},\lambda_{m})$, we use optimization (the Python scipy.optimize.minimize function) to find a current configuration, $\{x_{1},x_{2},x_{3}\}$, which realizes a particular optical switch for each wavelength and each input port. Numerical optimization was used to minimize the infidelity function Inf=1FInf=1-F with fidelity defined as
@@
Finally, the optimized current configurations were set on the photonic chip, and the corresponding optical switching was observed. Figure4c shows the average port-to-port switching fidelity among all switching configurations for each wavelength of coherent radiation. It can be seen that while the absolute values of the average port-to-port switching fidelities are not ideal (all less than 0.98) for each wavelength, the mutual average fidelity between the measured output power distribution and its prediction by a digital model is at least 0.99 (see Appendix F for details). This clearly shows that if the target interferometer is unable to implement a required transfer matrix, the digital model provides an accurate estimation of the expected fidelity.
It is worth noting that although the necessary currents for the realization of a particular unitary transformation $U(\vec{x}$ on the chip were obtained through optimization, this optimization process was not performed directly on the chip, as previously reported in [32]. Instead, it was conducted by using a digital chip model. Therefore, using a digital model effectively saves the resources of the target photonic chip and helps extend its active lifetime by avoiding the need for excessive reconfiguration. Additionally, in Appendix G we discuss the scalability of the proposed calibration method which in terms of the required single heater calibrations grows quadratically with the number of ports $N$ in the chip, and in Appendix H we describe an alternative optical chip calibration procedure based on the training of an architecture-agnostic model on 100 measured unitary matrices on the chip.

### v2

@@ Abstract
Reconfigurable photonics have rapidly become an invaluable tool for information processing. Light-based computing accelerators are promising for boosting neural network learning and inference [1, 2, 3], and optical interconnects are foreseen as a solution to the information transfer bottleneck in high-performance computing [4, 5, 6]. In this study, we demonstrate the successful programming of a transformation implemented using a reconfigurable photonic circuit with a non-conventional architecture. The core of most photonic processors is an MZI-based architecture [7, 8] that establishes an analytical connection between controllable parameters and circuit transformation. However, several architectures [9, 10] that are substantially more difficult to program have improved robustness to fabrication defects. We use two algorithms that rely on different initial datasets to reconstruct the circuit model of a complex interferometer, and then program the required unitary transformation. Both methods performed accurate circuit programming with an average fidelity greater than 98%. Our results provide a strong foundation for the introduction of non-conventional interferometric architectures for photonic information processing.
@@ I Introduction
Integrated photonics has significantly affected communications [11, 12, 13], quantum and classical computing [14], optical neural networks [15, 2, 3], and life sciences [16, 17]. Its core capability—on‑chip phase modulation—enables the routing and manipulation of optical states, allowing interferometric waveguide circuits to implement programmable information processing algorithms in light. These systems offer a wide operational bandwidth and superior energy efficiency per operation [18, 19, 20].
In this study, we focus on photonic device programming for information processing. A typical photonic chip for this purpose includes multiple power splitters and phase shifters comprising a large multiarm interferometer [21, 22, 23]. Once a photonic processor has been fabricated, it must be properly calibrated to implement a unitary transformation on demand. This process is challenging if the architecture of the interferometric circuit includes complex elements, such as multiport power splitting circuits [24, 2, 3].
@@
Common photonic architectures rely on a Mach-Zehnder interferometer (MZI) as a building block [7, 8, 25, 26]. This class of architectures is based on the well-known isomorphism of the MZI transfer matrix and arbitrary unitary rotation $SU(2)$. Within MZI-based architectures, calibration can be performed for each interferometer one after another, as it is possible to isolate the effect of each individual block [21, 15, 27, 28, 29, 19].
However, this strategy is not applicable to interferometers with other types of building blocks. For example, using multiport beamsplitters [10] or beamsplitter meshes [30] turns out to be beneficial in minimizing the effects of fabrication imperfections. Interferometers of this kind have already been demonstrated [9, 31, 32], but programming a desired unitary transformation remains an unresolved challenge. The well-known issue of the crosstalk between multiple tunable phase shifters complicates the unitary programming problem even more by rendering easy calibration procedures inaccurate even in basic architectures such as [7, 8].
@@
However, there exists substantial practical interest in interferometers with $n\times n$ waveguide splitters [36, 37, 38], as they can be used as a powerful tool for photonic convolutional networks and random matrix generation [2, 39]. Meanwhile, there is no straightforward approach to enable the programming of unitary transfer matrices of the interferometer with $n\times n$ splitters.
It should be mentioned that many of the existing studies on MZI-based interferometer programming have not used the actual matrix fidelity measure to demonstrate the ability to achieve the desired unitary matrix set on a chip. Instead, in these studies, another metric called amplitude fidelity was used [28, 19, 34, 23, 35], which is inherently insensitive to the angles of the unitary transformation that are crucial for the evolution of multiphoton states through linear optical interferometers. However, amplitude fidelity has been argued to contain nearly the same amount of information, demonstrating the universal capabilities of the interferometers under study [29].
Several recent theoretical proposals suggested ways to program an arbitrary complex interferometer [40, 41]. In these proposals, the digital model of the interferometer is inferred from the experimental data set and then serves as a reliable tool to determine the phase configuration corresponding to the required unitary transfer matrix. Both methods [40, 41] require precise control of phase shifts $P(\vec{\varphi})$ which, in turn, demand preliminary fine calibration $\vec{\varphi}=\vec{\varphi}(\vec{X})$ of the phase modulators.
@@ II.1 Optical Chip Design
We demonstrate our programming algorithm on a photonic chip that was manufactured using femtosecond laser writing (FSLW) in a fused silica glass sample [42]. The description of the waveguide fabrication regime can be found in the Appendix A. The single-mode waveguide structure of the chip contained a target $4\times 4$ multiport interferometer and a preparation interferometer. The preparation interferometer is a mesh of three MZIs which can set an arbitrary amplitude distribution at the output of the preparation part. We use the preparation interferometer as a tool for calibrating the target interferometer and probing its transfer matrix. Both the preparation and the target interferometers are equipped with tunable thermo-optical phase shifters, which are also fabricated using the femtosecond laser writing setup.
@@
Calibration of the MZIs-based interferometer is a known procedure that is performed by individually calibrating each of its constituent MZIs [21, 27, 28, 19]. We used the same principles to calibrate the preparation interferometer. More details on the optical chip design and preparation for interferometer calibration can be found in Appendices B and C.
@@ II.2 Target interferometer calibration
However, in most real optical chips, even those fabricated using the most advanced technology [28], there is often an effect of mutual cross-influence between different phase shifters. This can occur because of thermal or electrical crosstalk. Due to the crosstalk effect, the optical power $P_{m}$ of the $m$-th output port of the chip can no longer be described by 2 because the approximation constants $A_{m}$ and $B_{m}$ are functions of $x_{j}$, as well as of the indices of the input and output ports of the interferometer. This leads to a complex and non-periodic dependence $P_{m}(x_{j})$. Figure 2b shows the typical dependency $P_{m}(x_{j})$ in the presence of crosstalk.
@@
where $x_{j}$ is the current applied to the $j$-th heater, $\varphi^{(i)}_{0}$ is a constant bias phase shift that exists in the particular interferometer arm, and $\alpha_{ij}$ is a constant that determines the strength of the effect of the $j$-th heater on the $i$-th phase shifter. The relation between the phases and the currents flowing through the heaters can be rewritten in matrix form as follows:
@@
Moreover, any experimentally measured optical power $P_{km}$ from the applied current for any input $k$ and output $m$ ports of the interferometer and for any heater $j$ must obey (5) with the same interferometer parameters $M_{1,2}$, $\overrightarrow{\Phi_{0}}$ and $A$, as they are all generated by the same photonic chip, described by its unitary matrix $U$. Once all possible optical powers from the applied current dependencies, $P_{km}(x_{j})$, have been experimentally measured for each $k$, $m$, and $j$, they can be simultaneously approximated using a single interferometer model, determining the constituents of the optical chips $M_{1,2}$ and the necessary calibration relations between all currents and phase shifts $\overrightarrow{\Phi_{0}}$ and $A$. The key points of the approximation over the whole dataset are given in the next subsection.
@@ II.3 The target interferometer model
where $\vec{t}_{1,2}=\{t_{1},\dots,t_{N}\}$ are two sets of real parameters that determine $M_{1,2}$, $P(\vec{\varphi}(\vec{x},\vec{\varphi}_{0},A))=diag(e^{i\varphi_{1}},e^{i\varphi_{2}},e^{i\varphi_{3}},1)$ is a diagonal matrix of phase shifts (see Fig. 2c). In turn, phase shifts depend on the currents $x_{j}$ and the calibration parameters $\varphi^{(0)}_{j}$ and $\alpha_{ij}$:
@@
For mode-mixing blocks $M_{1,2}$ we used the well-known unitary parameterization [7] that for the $4\times 4$ port interferometer requires nine real parameters varying from 0 to $2\pi$ each (see Appendix D for details). The total number of real parameters for our photonic chip then sums up to $30$, including $9\times 2=18$ parameters for $M_{1,2}$, $3$ bias phases $\varphi^{(0)}_{j}$, and $9$ crosstalk constants $\alpha_{ij}$.
@@ III Experimental results
We gather the data set $P_{km}(x_{j})$ and run the approximation which yielded $M_{1}$ and $M_{2}$, $\overrightarrow{\Phi_{0}}$ and $A$, which are illustrated in Fig. 2d,e. Although no additional constraints were set for parameterization of $M_{1}$ and $M_{2}$, they appeared to be almost identical $M_{1}\approx M_{2}$ as shown in Fig. 2d. This fact comes from the stability of our FSLW setup. Moreover, from Fig. 2e we can see that the bias phases satisfy $\varphi^{0}_{1}\approx\varphi^{0}_{4}$ and $\varphi^{0}_{2}\approx\varphi^{0}_{3}$, which can also be explained by the actual geometry of the phase shift layer $P(\varphi)$ on our chip (see Fig. 1). In the phase-shift layer of the target interferometer, the first and fourth waveguides have the same length, while the second and third waveguides also have the same length but are shorter than the first and fourth. This results in non-zero bias phase shifts, as shown in Fig. 2e.
The reconstructed elements of the crosstalk matrix $A$ (see Fig. 2e) correlate with our assumption about their mutual relation based on the geometry of the optical structure. The negative values of some of the elements are also predicted by our simple model. Qualitative estimation of relative magnitudes of $a_{ij}$ can be made using the linearity of the heat equation, the assumption of logarithmic temperature decay with distance from the heater and the proportionality between phase and temperature differences $\varphi_{i}-\varphi_{j}\sim T_{i}-T_{j}$. Assuming (endless) parallel straight heaters with a width of $r_{0}$ right above (endless) parallel straight waveguides are equally separated by $d$ from each other (see Fig. 3), one can derive the following form of the crosstalk matrix $A$:
@@ III.1 Testing the digital model of the chip
To verify the validity of our digital interferometer model, we performed the following procedure. We sampled a set of 100 random current configurations $\{x^{(n)}_{1},x^{(n)}_{2},x^{(n)}_{3}\}$ ($n=\overline{1,100}$) in the range of 0 to 15 mA and applied them to our photonic chip. For each set of currents, the unitary transformation of the chip $U^{(n)}_{exp}$ was measured directly using coherent CW radiation [43, 44] (see Appendix E for details). On the other hand, $U^{(n)}$ can be simulated from the digital chip model (6) and compared to the one measured by the matrix fidelity measure:
@@
III.2 Machine learning based method for digital model reconstruction
We performed a proof-of-concept experiment to reconstruct the digital interferometer model and perform transfer matrix programming according to [40], which is based on the different data sets.
This method uses the training and validation data sets comprised of pre-reconstructed unitaries corresponding to the values of the driving signals $\vec{x}$ (currents through the heaters in our case) or the phase shift values $\vec{\varphi}$ induced by the driving signals. We used the calibration $\varphi(x)$ and the unitary matrices that were previously measured (see Fig. 4a) and prepared two data sets. The first data set $(\vec{\varphi},U)$ includes the phase shifts $\vec{\varphi}$ and the corresponding matrices $U$ and the second data set $(\vec{x},U)$ includes the current values $\vec{x}$ and the corresponding matrices $U$. We then used both data sets to estimate the performance of the method and compare it with the results provided by the global optimization. The details of the learning process can be found in Appendices F and G.
The method [40] infers the digital interferometer model
$U=A_{2}(\vec{p}_{2})\times P\times A_{1}(\vec{p}_{1}),$ (10)
where $P$ is a diagonal matrix of phase shifts and $A_{1},A_{2}$ are complex matrices, parameterized with $\vec{p}_{1},\vec{p}_{2}$ variables correspondingly. The diagonal matrix $P$ includes phase shifts $\vec{\varphi}$ or current-dependent phase functions $e^{i\varphi_{i}(x)}$ (see Appendices F and G for details). Data sets $(\vec{\varphi},U)$ and $(\vec{x},U)$ were divided into a training and a validation set. The training set was used to learn $A_{1,2}$ and $P(\vec{x)}$. The interferometer model was then evaluated using the validation set. The results are presented in Fig. 5a. As can be seen from the results, the fidelity between the measured unitary matrices and the predicted by both $(\vec{\varphi},U)$ and $(\vec{x},U)$ models is no less than $0.90$ with a peak around $0.99$, which indicates the precision of the reconstructed digital interferometer model using the method of [40].
However, the phase shifts-matrices dataset requires knowledge of the phase shift information, which is obtained from the calibration procedure that provides the mode mixing elements $M_{1,2}$ by itself. Therefore, performing another optimization procedure to determine the $A_{1,2}$ matrices, while already having the $M_{1,2}$ may seem somewhat redundant. In this regard, a more fundamental dataset of currents and corresponding unitary matrices for chips’ transformation is in high demand.
It is worth mentioning that in this experiment, training the artificial network on a currents-matrices dataset ($\vec{x}$ - $U$) to determine the ML-based optical chip model aims to solve exactly the same task as the calibration of the phase shift layer. Both methods return the digital chip model as a function of the experimentally tunable parameters: electrical currents $\vec{x}$. This model enables the prediction of the unitary transformation of the chip at a given current. However, despite sharing the same goal, these two methods differ in the experimental data they utilize. The calibration method requires all the output power from the current dependencies, whereas the ML method requires a set of unitary matrices measured on the chip. Under realistic experimental conditions, it may not always be easy to perform full optical chip tomography, in contrast to simple output power calibration measurements. At least, because if we use the same experimental resources (coherent laser radiation), full chip tomography would require injecting the radiation into two ports of the chip simultaneously and precisely varying their relative phase shift. However, only one input port was used in the calibration measurements. Then both methods use optimization over a set of real-valued parameters to minimize the corresponding loss function.
III.3 Optical port-to-port permutations in a wide range
@@
After a successful approximation of the entire data set of the calibration data collected, a complete digital model of the chip is obtained. In this procedure, the wavelength $\lambda$ determines the resulting unitary transformation $U=U(\lambda)$ due to the spectral features of the mode-mixing blocks and the spectral dependence of the phase shifts. Reconstruction of the digital model can be repeated with different wavelengths $\lambda_{new}$ to retrieve the chip model $U=U(\lambda_{new})$. The multiport interferometer architecture [10] was theoretically proven to be robust to the deviations of the mode-mixing block transfer matrices. By implementing the same operation, the optical port-to-port switching, using interferometers with different mode-mixing blocks, we highlight the robustness feature of the multiport architecture.
@@
To demonstrate the robustness of our photonic chip to the imperfections in the mode-mixing layers and its potential to perform predefined optical transformations with practical applications, we have experimentally performed optical switching at seven different wavelengths between $915$ and $975$ nanometers. The optical switching is a particularly demanding task for an interferometric circuit, especially based on MZI blocks, because it requires precise calibration of the power couplers and phase shifters. We first calibrated our optical chip for seven wavelengths using a continuously tunable CW laser (Toptica CTL 950), which has a range of wavelengths between $\lambda_{m}=915$ and $975$. We then obtained digital models $U(\vec{x},\lambda_{m})$ for each of these wavelengths. Then, using the known digital model of the chip, $U(\vec{x},\lambda_{m})$, we use optimization (the Python scipy.optimize.minimize function) to find a current configuration, $\{x_{1},x_{2},x_{3}\}$, which realizes a particular optical switch for each wavelength and each input port. Numerical optimization was used to minimize the infidelity function Inf=1FInf=1-F with fidelity defined as
@@
Finally, the optimized current configurations were set on the photonic chip, and the corresponding optical switching was observed. Figure4c shows the average port-to-port switching fidelity among all switching configurations for each wavelength of coherent radiation. It can be seen that while the absolute values of the average port-to-port switching fidelities are not ideal (all less than 0.98) for each wavelength, the mutual average fidelity between the measured output power distribution and its prediction by a digital model is at least 0.99 (see the Appendix H for details). This clearly shows that if the target interferometer is unable to implement a required transfer matrix, the digital model provides an accurate estimation of the expected fidelity.
IV Discussion
It should be noted that although the necessary currents for the realization of a particular unitary transformation $U(\vec{x}$ on the chip were obtained through optimization, this optimization process was not performed directly on the chip, as previously reported in [32]. Instead, it was conducted using a digital chip model. Therefore, using a digital model effectively saves resources of the target photonic chip and helps to extend its active lifetime by avoiding the need for excessive reconfiguration. Additionally, it can be shown that the scalability of the proposed calibration method increases quadratically with the number of required single heater calibrations on the optical chip (for details, see Appendix I). This is a typical scaling factor for reconfigurable multi-port optical chip calibration routines.